OpenStudy (anonymous):

a bunch of flowers contains 3 carnation, 3 roses and 5 sunflowers. two flowers are picked from the bunch without replacement. Find the probability for selecting: 2 roses

OpenStudy (anonymous):

@amistre64 please help me i need help badly

OpenStudy (amistre64):

how many flowers are there entotal?

OpenStudy (anonymous):

11

OpenStudy (amistre64):

and how many roses?

OpenStudy (anonymous):

3

OpenStudy (amistre64):

then our best estimate for picking one rose is 3 out of 11 3/11 since we dont replace stuff; that leaves 2 roses out of 10 flowers left

OpenStudy (anonymous):

3/11 second is 3/10 because we dont replace i cant get the final number by my self

OpenStudy (anonymous):

i mean 2/10

OpenStudy (amistre64):

the setup for P(rose1 and rose2) is: P(rose1) * P(rose 2) 3/11 * 1/5

OpenStudy (anonymous):

i dont know how to multiply fractions in the calculator 3*5 is 15 and 11*1 is 11 so 11/15?

OpenStudy (amistre64):

you just multiply across the tops, then across the bottom; then divide the tops by the bottoms 3*1 = tops 11*5 = bottoms tops/bottoms

OpenStudy (anonymous):

3/55

OpenStudy (amistre64):

if it wants an exact result, its prolly best to keep it in fraction form; if it says to apprximate by so many decimals; then this way is suitable

OpenStudy (amistre64):

yes, so in the calculator; do 3/55

OpenStudy (anonymous):

oh its fractions form will you help me get through like three more you can just tell me what to do and ill do it

OpenStudy (amistre64):

ive got a paying job to give some attention to today, but ill be hovering about

OpenStudy (anonymous):

the next one is same question different scenario the scenario is 1 carnation and 1 sunflower

OpenStudy (amistre64):

#carnation #sunflowers --------- * ---------- # flowers #flowers

OpenStudy (anonymous):

well terenz can help

OpenStudy (amistre64):

since without replacement, be sure to haave #flowers decrease as needed

terenzreignz (terenzreignz):

what

OpenStudy (anonymous):

3/11 and 5/10 or 1/2

OpenStudy (amistre64):

good

OpenStudy (anonymous):

thanks for your help @amistre64

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

okay so 1/2 and 3/11 multiply across top right @terenzreignz

terenzreignz (terenzreignz):

multiply the numerators and denominators... safest way to multiply fractions, but not always efficient.

OpenStudy (anonymous):

3/22

terenzreignz (terenzreignz):

I realised I haven't even read the question yet... derp :D

OpenStudy (anonymous):

lol

terenzreignz (terenzreignz):

So, question is, if you pick two flowers, what are the chances that you get one carnation and one sunflower?

OpenStudy (anonymous):

yes without replacement

terenzreignz (terenzreignz):

Well, actually, if you just consider 3/11 times 5/10 you're very specific that you pick the carnation first, and then the sunflower But what about picking a sunflower first, and the carnation afterwards? What are the chances of that?

OpenStudy (anonymous):

5/11 and 3/10 but it has carnation first lol

terenzreignz (terenzreignz):

oh, carnation first, specifically? Then forget what I said :)

OpenStudy (anonymous):

it says 1 carnation and 1 sunflower i assumed it wanted them in that order i may be wrong in assuming but

terenzreignz (terenzreignz):

but... please be sure, because if order doesn't matter, we should get an entirely different answer

OpenStudy (anonymous):

i dont know i guess it doesnt let me ask my teacher

terenzreignz (terenzreignz):

Please do :)

OpenStudy (anonymous):

he said order doesnt matter

terenzreignz (terenzreignz):

Ahah... then you should also consider the possibility that you pick a sunflower first, and the carnation afterward... could you figure that? :)

OpenStudy (anonymous):

so carnation first then sunflower is 3/11 and 5/10 then the sunflower first then the carnation is 5/11 and 3/10

terenzreignz (terenzreignz):

yup... so carnation first gives you 3/22, right? What about sunflower first?

OpenStudy (anonymous):

yeah sunflower first gives you : 15/110 i think and i need to simplify i think

terenzreignz (terenzreignz):

They're both divisible by 5, so, it gives you 3/ 22 as well... this is very interesting :D

OpenStudy (amistre64):

just so you know, multiplication is commutative

OpenStudy (anonymous):

yes which means the answer is 3/22 right?

OpenStudy (amistre64):

\[\frac{3}{11}\frac{5}{10}=\frac{5}{11}\frac{3}{10}\]

OpenStudy (anonymous):

whats that @amistre64

terenzreignz (terenzreignz):

crud i missed that detail... I've been away from here for far too long :D

OpenStudy (amistre64):

commutative means that numbers can switch places ... for example: 3 * 7 is equal to 21 swapping places doesnt change the value of the operation 7 * 3 = 21

OpenStudy (anonymous):

yes so both sides are = to 3/22

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so the answer is just 3/22

terenzreignz (terenzreignz):

lol no wonder... I can't believe I even marvelled at that :D so anyway, the chances of drawing a sunflower and a carnation, without respect to order, would be the sum of the chances of drawing a sunflower first, and then a carnation, and the chances of drawing a carnation first, and then a sunflower, both of which are 3/22

OpenStudy (anonymous):

lol marveled such a big descriptive word why do i feel like i am at a debate competition all of a sudden

terenzreignz (terenzreignz):

No, I completely submit :D Anyway, chances of drawing a sunflower first and then a carnation is 3/22 chances of drawing a carnation first and then a sunflower is (also) 3/22 Add those up, what do you get?

OpenStudy (anonymous):

6/44 which simplifies to 3/22

terenzreignz (terenzreignz):

oh...uhh No Big no. In adding fractions with the same denominator you add ONLY the numerators and NOT the denominators. Let me demonstrate... two halves make a whole, right? But if you also add denominators, then... 1/2 + 1/2 = 2/4 which simplifies into 1/2 implying 1/2 + 1/2 = 1/2 two halves make a half? No way :D So yeah, I hope I've made my point XD

OpenStudy (anonymous):

okay so its 6/22

terenzreignz (terenzreignz):

Yup... which simplifies into...?

OpenStudy (anonymous):

crud lol ughm 3/11

terenzreignz (terenzreignz):

That's the one :D

OpenStudy (anonymous):

okay let me make it a new question but i have a nother

OpenStudy (anonymous):

ill tag you

terenzreignz (terenzreignz):

Will do :)